y'=(4x+3y+15)/(2x+y+7)


1
Expert's answer
2021-02-24T06:31:54-0500

"\frac{dy}{4x+3y+15}=\frac{−dx}{2x+y+7}=\frac{dy−λdx}{4x+3y+15+λ(2x+y+7)}"

"=\frac{d(y−λx)}{(λ+3)y+(2λ+4)x+7λ+15}"

"=\frac{d((λ+3)y−λ(λ+3)x)}{(λ+3){(λ+3)y+(2λ+4)x+7λ+15}}"

For consistency in numerator and denominator

"\u2212\u03bb(\u03bb+3)=2\u03bb+4\u27f9"

"\u03bb2+5\u03bb+4=0\u27f9\u03bb=\u22121or\u22124"

Equating two ratios with the two "\u03bb \u2019s"

"\frac{d((−1+3)y−(−1)(−1+3)x)}{(−1+3){(−1+3)y+(2⋅−1+4)x+7⋅−1+15}}=\frac{d((−4+3)y−(−4)(−4+3)x)}{(−4+3){(−4+3)y+(2⋅−4+4)x+7⋅−4+15}}"

"\frac{d(2y+2x+8)}{2{2y+2x+8}}=\frac{d(−y−4x−13)}{−1{−y−4x−13}}"

"\frac{d(y+x+4)}{{y+x+4}}+2\frac{d(y+4x+13)}{{y+4x+13}}=0"

Integrating,

"ln(y+x+4)+2ln(y+4x+13)+lnC"

"(y+x+4)(y+4x+13)2=C"

Verification

since my answer is at variance to other answers it needs verification.

Differentiating (1)

"2(y+x+4)(dy+4dx)+(y+4x+13)(dy+dx)=0"

"(2y+2x+8+y+4x+30dy+(8y+8x+32++y+4x+13)dx=0"

"(6x+3y+21)dy+(12x+9y+45)dx=0"

"(2x+y+7)dy+(4x+3y+15)dx=0"

Verified.